OpenStudy (anonymous):

Find B^-1 Matrice: [-3 -1] [-4 -2]

hero (hero):

Which method do you have to use to find the inverse?

OpenStudy (anonymous):

Nothing specifically. @Hero

hero (hero):

Are you aware of the methods that currently exist to find the inverse of a matrix?

OpenStudy (anonymous):

I'm just checking this for a friend, I'm not sure how to do this anymore lol.

hero (hero):

The two most commonly used computational methods are: 1. Augmented Matrix method 2. Adjoint method As detailed in this link: http://www.mathwords.com/i/inverse_of_a_matrix.htm

OpenStudy (anonymous):

Can you show me how to do that?

hero (hero):

Basically \[B^{-1} = \frac{1}{\det(B)}\begin{bmatrix} d & -b \\ -c & a \end{bmatrix}\]

hero (hero):

So first find the determinant of B which is \(ad - bc\) or rather \((-3)(-2) -(-4)(-1)\)

hero (hero):

so \(\det(B) = 6 - 4 = 2\)

OpenStudy (anonymous):

So detB is = 2?

hero (hero):

Yes, correct. Now, swap \(a\) and \(d\) and then make \(c\) and \(b\) positive to get:

OpenStudy (anonymous):

Okay so 1/2 [-2 1] [4 -2]

hero (hero):

\[B^{-1} = \frac{1}{2}\begin{bmatrix} -2 & 1 \\ 4 & -3 \end{bmatrix}\]

OpenStudy (anonymous):

Oh whoops missed the -3. Now do I muliply each or what?

hero (hero):

Now you do this: \[B^{-1} = \begin{bmatrix} -2/2 & 1/2 \\ 4/2 & -3/2 \end{bmatrix}\]

hero (hero):

Basically divide each value inside the matrix by half

hero (hero):

Which afterwards you should end up with: \\[B^{-1} = \begin{bmatrix} -1 & \frac{1}{2} \\ 2 & -\frac{3}{2} \end{bmatrix}\]

OpenStudy (anonymous):

Awesome that's what I got, thanks so much!